OpenStudy (anonymous):

If ON=5x-5, LM=4x+4, NM=x-9 and OL=2y-5. What's the value of x and y for which LMNO must be a parallelogram ?????

OpenStudy (anonymous):

 set 4x+4 = 5x -5 solve for x, which is 9 then plug in y in terms of x. x-9=2y-5 so (9) -9 = 2y-5 solve for y, which is 2.5 remember to put the answer in units! :)

OpenStudy (anonymous):

So what is x, and y is 2.5?

OpenStudy (anonymous):

x is 9 y is 2.5